the ratio of dexter's salary to claudia's 4 to 5. if claudia earns $3200 per month. how much of a raise will dexter have to receive to make the ratio of his salary to claudia's 5 to 6?
Answer by math_tutor2020(3816)


d = dexter's monthly salary
c = claudia's monthly salary = 3200

d/c = 4/5
d = (4/5)*c
d = (4/5)*3200
d = 2560

r = raise amount

(d+r)/c = 5/6
(2560+r)/3200 = 5/6
6(2560+r) = 3200*5
15360+6r = 16000
6r = 16000-15360
6r = 640
r = 640/6
r = 106.67 approximately

Answer: $106.67